# Hadwiger Models: Low-Temperature Behavior in a Natural Extension of the Ising Model

### v1

@@ 3.3 Non-Peierls Lines
With reflection positivity we are able to use the chessboard estimate. Define a “block” as the fundamental domain of the set of reflections including every horizontal reflection and every other vertical reflection, displayed in Figure 8. This will include one vertex in its center, and two along its edge. For a local event $A$ on a single block $B$, we can define an event $A_{i}$ on any other block $B_{i}$ by applying the necessary reflections to map that block to $B$. Notate by $\alpha$ the global event that $A_{i}$ occurs simultaneously on all blocks, indexed by $S$. The chessboard estimate states [FV17]:
$\mathbb{P}(A)\leq\mathbb{P}(\alpha)^{1/|S|}\,.$
Theorem 7.
Along the $E-C$ transition line, if $e_{F}\geq e_{H}$, the probability a vertex is in the $H$ or $F$ states decays exponentially in the inverse temperature
Proof.
We decompose the event into $\sigma=F$ and $\sigma=H$. First, consider $\sigma=F$. In this case, if $\sigma_{i}=F\forall i\in S$, then every hexagon is filled, so $\sigma_{i}=F,\forall i$, which corresponds to exactly 1 configuration. The total number of vertices is twice the number of blocks, so the weight of this configuration is $e^{-2e_{F}|S|/T}$ and the partition function of the whole system is bounded below by 1, so $\mathbb{P}(\sigma=F)<(e^{-2e_{F}|S|/T})^{1/|S|}=e^{-2e_{F}/T}$. Now we consider $\sigma=H$. There are no more than $3^{|S|}$ configurations such that $\sigma_{i}=H,\,\forall i\in S$, because each block can take no more than 3 independent configurations with state $S$. The non-central vertices may be in a lower energy state, but the energy of the total configuration must be at least $e_{H}|S|$, so $\mathbb{P}(\sigma=H)<(3^{|S|}e^{-e_{F}|S|/T})^{1/|S|}=3e^{-e_{F}/T}$. Thus $\mathbb{P}(\sigma=H\text{ or }F)<e^{-2e_{F}/T}+3e^{-e_{F}/T}$. The same statement holds along the $H-F$ transition line, when $e_{E}\geq e_{C}$. ∎
With this, we can make the connection to the hard hexagon model explicit at sufficiently low temperatures. For a fixed infinite configuration $\eta$, notate the probability of an event $A$ for the subcritical (i.e. no domination) Hadwiger model at temperature $T$ on the finite region $R$ with boundary conditions induced by $\eta$ as $\mathbb{P}^{R}_{S,\eta,T}(A)$ and the distribution of the Hard Hexagon model with $z=1$ (i.e. even distribution) on the same region with the same boundary conditions by $\mathbb{P}^{R}_{H,\eta}$. Notate their respective infinite volume limits by $\mathbb{P}_{S,\eta,T}(A)$ and $\mathbb{P}_{H,\eta}$.
@@
$|\mathbb{P}^{M}_{S,\eta_{1},T}(A)-\mathbb{P}^{M}_{S,\eta_{2},T}(A)|\leq\mathbb{P}_{p_{i}}(\text{there is an open path from some vertex in }N\text{ to }\partial M)$
where $\mathbb{P}_{p_{i}}$ is Bernoulli site percolation with probability $p_{i}$, and $p_{i}$ is the variational distance between neighbor boundary defined in the theorem on (Disagreement Percolation).. $p_{i}<1/2$ at any non-zero temperature, so the probability of an open path is always less than at the critical probability, $1/2$. For fixed $N$, we can choose $M_{0}$ such that [K82]
$\mathbb{P}_{1/2}(\text{there is an open path from some vertex in }N\text{ to }\partial M)<\epsilon/2$
Notate by $A^{M}_{EC}$ the event that no vertex in $M$ is in the state $F$ or $H$. Given $M_{0}$, we can choose $T$ such that $\mathbb{P}^{M}_{S,\eta,T_{0}}(A^{M}_{EC})>1-\epsilon/2$, by the previous theorem. $\mathbb{P}^{M}_{H,\eta}(A)=\mathbb{P}^{M}_{S,\eta,T}(A|A^{M}_{EC})$, for any $T$, by the definition of the hard hexagon model, so
@@
$\min_{\eta^{\prime}}\mathbb{P}^{M}_{S,\eta^{\prime},T}(A)\leq\mathbb{P}^{M^{\prime}}_{S,\eta,T}(A)\leq\max_{\eta^{\prime}}\mathbb{P}^{M}_{S,\eta^{\prime},T}(A),\text{ and }\min_{\eta^{\prime}}\mathbb{P}^{M}_{H,\eta^{\prime}}(A)\leq\mathbb{P}^{M^{\prime}}_{H,\eta}(A)\leq\max_{\eta^{\prime}}\mathbb{P}^{M}_{H,\eta^{\prime}}(A)$
@@ 4 Special Cases
To our knowledge, the only finite-energy model in the Hadwiger class other than the Ising model in a field that has been exactly solved is the triplet model, sometimes called the Ising model with three-spin interactions [B74]. This has Hamiltonian $-\sum_{v}h_{v1}h_{v2}h_{v3}$, where $h_{v1},h_{v2},h_{v3}$ are the spins in $\{-1,1\}$ of the hexagons adjacent to $v$, which results in a model at the midpoint of the $C-F$ or $E-H$ transition lines, depending on sign convention. This makes it a natural analogue of the Ising model in the configuration space, which is the midpoint of the only other Peierls transition line. The triplet model exhibits a sharp phase transition, which, given the sharp transitions in the Ising model as well, is reason to believe that in fact the entire space of Hadwiger functions experiences such a transition, except along the non-Peierls transition lines.
@@ 5 Further Directions
As any finite-range model displays only a single undominated Gibbs state at sufficiently high temperatures [FV17], models in the $C$ and $H$ regions must experience a change in the number of Gibbs state at some intermediate temperature. However, precisely where this change occurs, and if there are temperatures with a different number of Gibbs states, is yet unknown. For any set of parameters not on the non-Peierls transition lines, a transition occurs in the sense that some configurations dominate at low temperature, while at high temperatures no configuration dominates. However, the nature of this transition is also not known.
@@
In 3 dimensions, the space of Hadwiger functions is 4-dimensional: Euler characteristic, mean width, surface area, and volume [LMNPT20]. Here, “mean width” refers to the average projected length of the set, over the space of 1-dimensional subspaces. The natural analogue to the lattice of hexagons is the lattice of truncated octahedra; for both, $d+1$ identical cells meet at each vertex. However, unlike the hexagonal lattice, for which the Euler characteristic is antisymmetric, the Euler characteristic is now symmetric, paired with the symmetric surface area term. Volume remains antisymmetric, along with the new mean-width terms. Up to temperature normalization, this generates two 1-dimensional subspaces of purely symmetric or anti-symmetric energy functions, granting the model many more potential symmetries to rely on.
6 Acknowledgements
We’d like to thank Alexander Smith at the University of Minnesota as well as Érika Roldán and Ron Peled for essential conversations on the paper. In particular, Smith noticed the connection between the class of Hadwiger functions and the Ising Hamiltonian, which inspired the whole project.
[…]

### v2

@@ 3.3 Non-Peierls Lines
With reflection positivity we are able to use the chessboard estimate. Define a “block” as the fundamental domain of the set of reflections including every horizontal reflection and every other vertical reflection, displayed in Fig. 9. This will include one vertex in its center, and two along its edge. For a local event $A$ on a single block $B$, we can define an event $A_{i}$ on any other block $B_{i}$ by applying the necessary reflections to map that block to $B$. Denote by $\alpha$ the global event that $A_{i}$ occurs simultaneously on all blocks, indexed by $S$. The chessboard estimate states [FV17]: $\mathbb{P}(A)\leq\mathbb{P}(\alpha)^{1/|S|}\,.$
Proof of Theorem 7.
We decompose the event into $\sigma=F$ and $\sigma=H$. First, consider $\sigma=F$. In this case, if $\sigma_{i}=F\forall i\in S$, then every hexagon is filled, so $\sigma_{i}=F,\forall i$, which corresponds to exactly 1 configuration. The total number of vertices is twice the number of blocks, so the weight of this configuration is $e^{-2e_{F}|S|/T}$ and the partition function of the whole system is bounded below by 1, so
$\mathbb{P}(\sigma=F)<(e^{-2e_{F}|S|/T})^{1/|S|}=e^{-2e_{F}/T}.$
Now we consider $\sigma=H$. There are no more than $3^{|S|}$ configurations such that ${\sigma_{i}=H},\,\forall i\in S$, because each block can take no more than 3 independent configurations with state $S$. The non-central vertices may be in a lower energy state, but the energy of the total configuration must be at least $e_{H}|S|$, so
$\mathbb{P}(\sigma=H)<(3^{|S|}e^{-e_{F}|S|/T})^{1/|S|}=3e^{-e_{F}/T}.$
Thus $\mathbb{P}(\sigma=H\text{ or }F)<e^{-2e_{F}/T}+3e^{-e_{F}/T}$. The same statement holds along the HFH-F degeneracy line, when $e_{E}\geq e_{C}$. ∎
With this, we can make the connection to the hard hexagon model explicit at sufficiently low temperatures. For a fixed infinite configuration $\eta$, denote the probability of an event $A$ for the subcritical (i.e. no domination) Hadwiger model at temperature $T$ on the finite region $R$ with boundary conditions induced by $\eta$ as $\mathbb{P}^{R}_{S,\eta,T}(A)$ and the distribution of the Hard Hexagon model with $z=1$ (i.e. even distribution) on the same region with the same boundary conditions by $\mathbb{P}^{R}_{H,\eta}$. Denote their respective infinite volume limits by $\mathbb{P}_{S,\eta,T}(A)$ and $\mathbb{P}_{H,\eta}$.
@@
[…]
where $\mathbb{P}_{p_{i}}$ is Bernoulli site percolation with probability $p_{i}$, and $p_{i}$ is the variational distance between neighbor boundary defined in Theorem 8. $p_{i}<1/2$ at any non-zero temperature, so the probability of an open path is always less than at the critical probability, $1/2$. For fixed $N$, we can choose $M_{0}$ such that [K82]
[…]
Denote by $A^{M}_{EC}$ the event that no vertex in $M$ is in the state $F$ or $H$. Given $M_{0}$, we can choose $T$ such that $\mathbb{P}^{M}_{S,\eta,T_{0}}(A^{M}_{EC})>1-\epsilon/2$, by the previous theorem. $\mathbb{P}^{M}_{H,\eta}(A)=\mathbb{P}^{M}_{S,\eta,T}(A|A^{M}_{EC})$, for any $T$, by the definition of the hard hexagon model, so
@@
$\min_{\eta^{\prime}}\mathbb{P}^{M}_{S,\eta^{\prime},T}(A)\leq\mathbb{P}^{M^{\prime}}_{S,\eta,T}(A)\leq\max_{\eta^{\prime}}\mathbb{P}^{M}_{S,\eta^{\prime},T}(A)$
$\min_{\eta^{\prime}}\mathbb{P}^{M}_{H,\eta^{\prime}}(A)\leq\mathbb{P}^{M^{\prime}}_{H,\eta}(A)\leq\max_{\eta^{\prime}}\mathbb{P}^{M}_{H,\eta^{\prime}}(A)$
@@ 4 Special Cases
To our knowledge, the only finite-energy model in the Hadwiger class other than the Ising model in a field that has been exactly solved is the Baxter-Wu model, sometimes called the Ising model with three-spin interactions [BW73]. This has Hamiltonian $\sum_{v}h_{v1}h_{v2}h_{v3}$, where $h_{v1},h_{v2},h_{v3}$ are the spins in $\{-1,1\}$ of the hexagons adjacent to $v$, which results in a model at the midpoint of the CFC-F or EHE-H degeneracy lines, depending on sign convention. This makes it a natural analogue of the Ising model in the configuration space, which is the midpoint of the only other Peierls degeneracy line. The Baxter-Wu model exhibits a sharp phase transition, which, given the sharp transitions in the Ising model as well, is reason to believe that in fact all Hadwiger models exhibit sharp phase transitions separating the different regions of the diagram, except along the non-Peierls degeneracy lines.
@@ 5 Further Directions
As any finite-range model displays only a single undominated Gibbs state at sufficiently high temperatures [FV17], models in the $C$ and $H$ regions must experience a change in the number of Gibbs state at some intermediate temperature. However, precisely where this change occurs, and if there are temperatures with a different number of Gibbs states, is yet unknown. For any set of parameters not on the non-Peierls degeneracy lines, a transition occurs in the sense that some configurations dominate at low temperature, while at high temperatures no configuration dominates. However, the nature of this transition is also not known.
@@
In three dimensions, the space of invariant valuations is four-dimensional, being parametrized by Euler characteristic, mean width, surface area, and volume [G07]. Here, mean width refers to the average projected length of the set over the space of 1-dimensional subspaces. The natural analogue to the lattice of hexagons is the lattice of truncated octahedra; for both, $d+1$ identical cells meet at each vertex. However, unlike the hexagonal lattice, for which the Euler characteristic is antisymmetric, the Euler characteristic is now symmetric, paired with the symmetric surface area term. Volume remains antisymmetric, along with the new mean-width terms. Up to temperature normalization, this generates two 1-dimensional subspaces of purely symmetric or anti-symmetric energy functions, granting the model many more potential symmetries to rely on.
Acknowledgments and Declarations
Acknowledgments: We’d like to thank Alexander Smith at the University of Minnesota as well as Érika Roldán and Ron Peled for essential conversations on the paper. In particular, Smith noticed the connection between the class of invariant valuations and the Ising Hamiltonian, which inspired the whole project. We’d also like to thank the anonymous referees for their useful comments and suggestions.
Competing Interests: We have no competing interests to declare.
Data Availability Statement: No datasets were generated or analyzed during the current study.